OpenStudy (anonymous):

area bounded by x=1 x=3 y=0 y=1/x

OpenStudy (anonymous):

\[\int_1^3\frac{dx}{x}\]

OpenStudy (anonymous):

aka \(\ln(3)\) pretty much by definition, since \[\ln(x)=\int_a^x\frac{dt}{t}\]